FIND THE GENERAL SOLUTION

Need ALL 4. Differential Equations class

Problem 1 Find the general solution 1. r(t) = (-2)=(0) o 1 1 2. x(t) = (1 0 1 ) r(t) 1 1 0 3. x(t) = 1) 2(0) 4. x(t) = ( =
Answer #1

① The given system is El 2-2 2(t) = (-3 12 2014) 12 - 2 it can be written as 1-32 X114) A>(t) ra-2 X2(t)) W; (4) = f O xi(t)where D=d H (D?+50+4) Xi(t)=0 (+4)(041) X(t) = 0 » (+4JX1(t)=0 B+1)xH)-6 =) comple general solution of ean 9 is x4 (t) = geht2 (t) = 1 1 1 (t) 1 1 1 0 it can be written as Xct) Xit) 1, Ct) ) 0 1 1 1 0 1 la 30 1 10 23(t) wit) = xt) + x36) X24) 2017) +D = d at 02-0-2) X(t) =0 . Auxillary en is m2m-2=0 (M-2) (Mt1=0 > m=2,- general somn of egno is Gact) = ge 24 Get = x; (t) =3t et et -t 72(t) = Cie + X2(A) = Ge the cetteget put Xg(t) & X₂(t) in eah ③ we get Iz ct) 2ge? + (624(3)é =) d 13(1) (29e2twe get مترا Differentiate o z = r; tixa substitute d2 we get x = x+i(-ix; + x2) xj + a +iX2 subtract * from ☺ we get 20-X =